OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (3, 6) and (−2, 1)?

OpenStudy (anonymous):

Hang on

OpenStudy (anonymous):

right

OpenStudy (anonymous):

can u help me? @JCMMusic

OpenStudy (anonymous):

Sorry just hold on please for one more minute like 5 people are all asking me questions!!

OpenStudy (anonymous):

haha ok . I'm waiting

OpenStudy (anonymous):

y = 1x + 3 I'm pretty sure is the answer is this a possible answer choice on your test or worksheet?

OpenStudy (anonymous):

a) y + 6 = 1(x + 3) b) y + 1 = −1(x + 2) c) y − 6 = −1(x + 3) d) y − 6 = 1(x − 3)

OpenStudy (anonymous):

so the answer is a) right?

OpenStudy (anonymous):

Hang on let me check

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

No it isn't a

OpenStudy (anonymous):

so which one is the answer?

OpenStudy (anonymous):

It's D but I can't help with any more questions

OpenStudy (anonymous):

I have to go ttyl

OpenStudy (anonymous):

it's ok. thank u